# Minimax estimation of functionals in sparse vector model with correlated observations

### v1

@@ 6.3.1 Analysis of $\hat{\sigma}_{S}^{2}$
$\left|F(t/\sigma^{2})-F(2t/3\sigma^{2}))\right|\geq\frac{2}{\sqrt{2\pi}}e^{-\frac{5}{3}}\left(\sqrt{\frac{1}{3}}-\sqrt{\frac{2}{9}}\right)\geq 0.015.$
@@
$\hat{q}_{\alpha}:=\min\{u:\hat{F}_{\varepsilon^{2}}(u)\geq\alpha\}.$
@@
${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}(\hat{q}_{\alpha}^{2})\leq 1+\frac{8}{(1-\alpha)\sqrt{2\pi}}.$
@@
${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}[1-\hat{F}_{\varepsilon^{2}}(t)]={\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}(\tilde{\varepsilon}_{1}^{2}>t)\leq\frac{2e^{-\frac{t}{2}}}{\sqrt{2\pi t}}.$
@@
[…]
@@
$\mathcal{T}:=\left\{t=2^{\ell}:\ell\in\mathbb{Z},\frac{\sigma^{2}}{3}\leq t\leq\frac{3}{2}\sigma^{2}\right\}.$
@@
${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\frac{1}{2}\sigma^{2}\leq\hat{\sigma}_{S}^{2}\leq\frac{3}{2}\sigma^{2}\right]\geq 1-c\frac{\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}^{2}}{d^{2}}.$
@@
Recall that, by definition, $\hat{\sigma}_{S}^{2}=\hat{\sigma}_{S}^{2}(\hat{t})$. We have
@@
${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left|\frac{\hat{\sigma}_{S}^{2}}{\sigma^{2}}-1\right|={\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left(\left|\frac{\hat{\sigma}_{S}^{2}}{\sigma^{2}}-1\right|\mathbbm{1}(E)\right)+{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left(\left|\frac{\hat{\sigma}_{S}^{2}}{\sigma^{2}}-1\right|\mathbbm{1}(E^{c})\right).$
@@
[…]
@@ 6.3.2 Analysis of $\hat{\sigma}^{2}_{D}$
${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\left(\frac{\hat{\varphi}_{d}(z)}{e^{-\frac{z^{2}\sigma^{2}}{2}}}-1\right)^{2}\right]\leq 2{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\left(\frac{\hat{\varphi}_{d}(z)-{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\hat{\varphi}_{d}(z)\right]}{e^{-\frac{z^{2}\sigma^{2}}{2}}}\right)^{2}\right]+2\left(\frac{2s}{d}\right)^{2}.$
@@
${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[|\hat{\varphi}_{d}(z)-{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}[\hat{\varphi}_{d}(z)]|\leq e^{-\frac{z^{2}\sigma^{2}}{2}}/4\right]\geq 1-16e^{z^{2}\sigma^{2}}\mathbf{Var}[\hat{\varphi}_{d}(z)].$
@@
$\tilde{\sigma}_{D}^{2}=-\frac{2\hat{t}}{\lambda}\log|\hat{\varphi}_{d}((\lambda/\hat{t})^{1/2})|$
@@
${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\frac{1}{2}e^{-\frac{\lambda\sigma^{2}}{2t}}\leq\hat{\varphi}_{d}((\lambda/t)^{1/2})\leq\frac{3}{2}e^{-\frac{\lambda\sigma^{2}}{2t}},\forall t\in\mathcal{T}\right]\geq 1-C\frac{\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}^{2}}{d^{2}}\left(\frac{s}{\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}}\vee 1\right)\exp\left(3\lambda\right).$
@@ 6.4 Proofs for Section 4 (unknown $\sigma$)
In this section, we prove Theorem 7 using the estimators of the noise level $\sigma^{2}$ constructed in Section 3. Under Assumptions LABEL:as:var0 and LABEL:as:var1, we have ${\boldsymbol{\Sigma}}_{ij}/(\sigma_{i}\sigma_{j})\asymp{\boldsymbol{\Sigma}}_{ij}$ and thus $\|{\boldsymbol{\Sigma}}\|_{F}\asymp\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}$. Hence, we can replace $\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}$ by $\|{\boldsymbol{\Sigma}}\|_{F}$ in the theorems of Section 3 and the bounds on estimation errors remain valid to within a constant factor.
@@ 6.4.1 Proof of part (i) of Theorem 7
$\left|\sqrt{\left|\|{\boldsymbol{\theta}}\|_{2}^{2}+2\sigma\boldsymbol{\theta}^{\top}{\boldsymbol{\varepsilon}}\right|}-\|{\boldsymbol{\theta}}\|_{2}\right|\leq\frac{2\sigma|\boldsymbol{\theta}^{\top}{\boldsymbol{\varepsilon}}|}{\|\boldsymbol{\theta}\|_{2}}.$
@@
$\left|\sum_{i=1}^{d}\Big{(}\varepsilon_{i}^{2}-\frac{\hat{\sigma}_{D}^{2}}{\sigma^{2}}\sigma_{i}^{2}\Big{)}\right|\leq\left|\sum_{i=1}^{d}(\varepsilon_{i}^{2}-\sigma_{i}^{2})\right|+\left|\sum_{i=1}^{d}\sigma_{i}^{2}\left(1-\frac{\hat{\sigma}_{D}^{2}}{\sigma^{2}}\right)\right|.$
Assumption LABEL:as:var0 implies that $\sum_{i=1}^{d}\sigma_{i}^{2}\leq d$. Therefore, we can use Theorem 6 to get
@@ 6.4.2 Proof of parts (ii) and (iii) of Theorem 7
Note that since $\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}\leq d$ it follows from Theorem 5 that ${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}(\hat{\sigma}_{S}^{2})\lesssim\sigma^{2}$. Moreover, Assumption LABEL:as:var0 implies that $\sum_{i\in\mathcal{S}}\sigma_{i}^{2}\leq s$. It follows that
@@
where (a) follows from Assumption LABEL:as:var0 and Theorem 5 while (b) is due to the second inequality in Proposition 2(ii), the assumption that $s\leq\|{\boldsymbol{\Sigma}}\|_{F}$, the inequality $\alpha_{s}\lesssim\tau^{2}e^{-\tau^{2}/2}$ and the fact that $\tau=3\sqrt{\log(1+\|{\boldsymbol{\Sigma}}\|_{F}^{2}/s^{2})}$.
@@
$(s+\|{\boldsymbol{\Sigma}}\|_{F})\tau^{3}e^{-\tau^{2}/4}\lesssim s\log(1+\|{\boldsymbol{\Sigma}}\|_{F}^{2}/s^{2})=\psi^{*}(s,{\boldsymbol{\Sigma}}).$
@@
$\mathcal{A}:=\left\{\frac{|\hat{\sigma}_{S}^{2}-\sigma^{2}|}{\sigma^{2}}\leq\delta:=\frac{c(s+\|{\boldsymbol{\Sigma}}\|_{F})}{2\eta d}\right\},$
@@
[…]
Consequently, since $\sigma_{i}^{2}\leq 1$ by Assumption LABEL:as:var0, we have
${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left(\Big{|}\sum_{i\not\in\mathcal{S}}\varepsilon_{i}^{2}\left(\mathbbm{1}\left(|\varepsilon_{i}|>\frac{\hat{\sigma}_{S}}{\sigma}\sigma_{i}\tau\right)-\mathbbm{1}\left(|\varepsilon_{i}|>\sigma_{i}\tau\right)\right)\Big{|}\mathbbm{1}(\mathcal{A})\right)\lesssim\frac{s+\|{\boldsymbol{\Sigma}}\|_{F}}{\eta}\tau^{3}e^{-\tau^{2}/4}.$
@@
${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\left\{\left|\sum_{i\not\in\mathcal{S}}\varepsilon_{i}^{2}\left(\mathbbm{1}\left(|\varepsilon_{i}|>\frac{\hat{\sigma}_{S}}{\sigma}\sigma_{i}\tau\right)-\mathbbm{1}\left(|\varepsilon_{i}|>\sigma_{i}\tau\right)\right)\right|\geq\frac{C(s+\|{\boldsymbol{\Sigma}}\|_{F})}{\eta^{2}}\tau^{3}e^{-\tau^{2}/4}\right\}\cap\mathcal{A}\right]\leq\eta/2,$
@@
where, by the Schur complement (see, e.g., [10, (0.7.3.1)]),
@@
${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\frac{\hat{\sigma}_{S}(t)}{\sigma}\wedge 1\leq\frac{|\varepsilon_{i}|}{\sigma_{i}\tau}\leq\frac{\hat{\sigma}_{S}(t)}{\sigma}\vee 1\right]\lesssim\sqrt{({\boldsymbol{\Sigma}}^{-1})_{ii}}\frac{s+\|{\boldsymbol{\Sigma}}\|_{F}}{d}e^{-\frac{\tau^{2}}{16}}+\left(\frac{s+\|{\boldsymbol{\Sigma}}\|_{F}}{d}\right)^{2}.$
@@
$\hat{\sigma}_{-i}^{2}(t):=\frac{t}{F^{-1}(\hat{F}_{-i}(t))},\qquad\delta_{-i}(t):=\left|\frac{\hat{\sigma}_{-i}^{2}(t)}{\sigma^{2}}-1\right|,$
@@
$\left|\frac{\hat{\sigma}_{-i}^{2}(t)}{\sigma^{2}}-\frac{\hat{\sigma}_{S}^{2}(t)}{\sigma^{2}}\right|\leq c/d,$
@@
$\iota_{i}^{2}=\frac{1}{\sigma_{i}^{2}({\boldsymbol{\Sigma}}^{-1})_{ii}}\in[0,1],\quad\mu_{i}({\boldsymbol{\varepsilon}}_{-i})\sim\mathcal{N}(0,1-\iota_{i}^{2}).$
@@
${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[|\varepsilon_{i}|/\sigma_{i}\in[l,u],E_{1}\right]\leq\leavevmode\nobreak$ ${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\mathbbm{1}(E_{1})\sqrt{\frac{2}{\pi}}\frac{L}{\iota_{i}}\exp\left(-\frac{(l-|\mu_{i}({\boldsymbol{\varepsilon}}_{-i})|)_{+}^{2}}{2\iota_{i}^{2}}\right)\right],$ (146)
@@
[…]
@@
[…]
@@
${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[L^{2}\mathbbm{1}(E_{1})\right]\lesssim\left(\frac{s+\|{\boldsymbol{\Sigma}}\|_{F}}{d}\right)^{2}.$
@@
${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\frac{\hat{\sigma}_{S}^{2}(t)}{\sigma^{2}}\wedge 1\leq\frac{\varepsilon_{i}^{2}}{\sigma_{i}^{2}\tau^{2}}\leq\frac{\hat{\sigma}_{S}^{2}(t)}{\sigma^{2}}\vee 1,E_{1}\right]\lesssim\sigma_{i}\sqrt{({\boldsymbol{\Sigma}}^{-1})_{ii}}\frac{s+\|{\boldsymbol{\Sigma}}\|_{F}}{d}\exp\left(-\frac{\tau^{2}}{16}\right).$
@@
${\rm E}_{1}=\{\hat{t}\in\mathcal{T}\},\quad{\rm E}_{2}=\left\{\hat{\sigma}_{S}^{2}\leq\frac{3}{2}\sigma^{2}\right\},$

### v2

@@ 6.3.1 Analysis of $\hat{\sigma}_{S}^{2}$
$\left|F(t/\sigma^{2})-F(2t/3\sigma^{2}))\right|\geq\frac{2}{\sqrt{2\pi}}e^{-\frac{5}{3}}\left(\sqrt{\frac{1}{3}}-\sqrt{\frac{2}{9}}\right)\geq 0.015.$ (118)
@@
$\hat{q}_{\alpha}:=\min\{u:\hat{F}_{\varepsilon^{2}}(u)\geq\alpha\}.$ (119)
@@
${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}(\hat{q}_{\alpha}^{2})\leq 1+\frac{8}{(1-\alpha)\sqrt{2\pi}}.$ (120)
@@
${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}[1-\hat{F}_{\varepsilon^{2}}(t)]={\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}(\tilde{\varepsilon}_{1}^{2}>t)\leq\frac{2e^{-\frac{t}{2}}}{\sqrt{2\pi t}}.$ (121)
@@
${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}[\hat{q}_{\alpha}^{2}]\leq 1+\int_{1}^{\infty}{\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}(\hat{q}_{\alpha}^{2}>t)dt\leq 1+\int_{1}^{\infty}\frac{2e^{-\frac{\sqrt{t}}{2}}}{(1-\alpha)\sqrt{2\pi}t^{1/4}}dt\leq 1+\frac{8}{(1-\alpha)\sqrt{2\pi}}.$ (123)
@@
$\mathcal{T}:=\left\{t=2^{\ell}:\ell\in\mathbb{Z},\frac{\sigma^{2}}{3}\leq t\leq\frac{3}{2}\sigma^{2}\right\}.$ (129)
@@
${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\frac{1}{2}\sigma^{2}\leq\hat{\sigma}_{S}^{2}\leq\frac{3}{2}\sigma^{2}\right]\geq 1-c\frac{\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}^{2}}{d^{2}}.$ (130)
@@
[…]
@@
${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left|\frac{\hat{\sigma}_{S}^{2}}{\sigma^{2}}-1\right|={\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left(\left|\frac{\hat{\sigma}_{S}^{2}}{\sigma^{2}}-1\right|\mathbbm{1}(E)\right)+{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left(\left|\frac{\hat{\sigma}_{S}^{2}}{\sigma^{2}}-1\right|\mathbbm{1}(E^{c})\right).$ (133)
@@
$\phantom{{}={}}{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\left|\frac{\hat{\sigma}_{S}^{2}}{\sigma^{2}}-1\right|\mathbbm{1}(E)\right]={\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\frac{\hat{\sigma}_{S}^{2}}{\hat{t}}\left|\frac{\hat{t}}{\sigma^{2}}-\frac{\hat{t}}{\hat{\sigma}_{S}^{2}}\right|\mathbbm{1}(E)\right]$ (137)
@@ 6.3.2 Analysis of $\hat{\sigma}^{2}_{D}$
${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\left(\frac{\hat{\varphi}_{d}(z)}{e^{-\frac{z^{2}\sigma^{2}}{2}}}-1\right)^{2}\right]\leq 2{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\left(\frac{\hat{\varphi}_{d}(z)-{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\hat{\varphi}_{d}(z)\right]}{e^{-\frac{z^{2}\sigma^{2}}{2}}}\right)^{2}\right]+2\left(\frac{2s}{d}\right)^{2}.$ (143)
@@
${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[|\hat{\varphi}_{d}(z)-{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}[\hat{\varphi}_{d}(z)]|\leq e^{-\frac{z^{2}\sigma^{2}}{2}}/4\right]\geq 1-16e^{z^{2}\sigma^{2}}\mathbf{Var}[\hat{\varphi}_{d}(z)].$ (145)
@@
$\tilde{\sigma}_{D}^{2}=-\frac{2\hat{t}}{\lambda}\log|\hat{\varphi}_{d}((\lambda/\hat{t})^{1/2})|$ (146)
@@
${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\frac{1}{2}e^{-\frac{\lambda\sigma^{2}}{2t}}\leq\hat{\varphi}_{d}((\lambda/t)^{1/2})\leq\frac{3}{2}e^{-\frac{\lambda\sigma^{2}}{2t}},\forall t\in\mathcal{T}\right]\geq 1-C\frac{\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}^{2}}{d^{2}}\left(\frac{s}{\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}}\vee 1\right)\exp\left(3\lambda\right).$ (154)
@@ 6.4 Proofs for Section 4 (unknown $\sigma$)
In this section, we prove Theorem 6 using the estimators of the noise level $\sigma^{2}$ constructed in Section 3. Under Assumptions 1 and 2, we have ${\boldsymbol{\Sigma}}_{ij}/(\sigma_{i}\sigma_{j})\asymp{\boldsymbol{\Sigma}}_{ij}$ and thus $\|{\boldsymbol{\Sigma}}\|_{F}\asymp\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}$. Hence, we can replace $\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}$ by $\|{\boldsymbol{\Sigma}}\|_{F}$ in the theorems of Section 3 and the bounds on estimation errors remain valid to within a constant factor.
@@ 6.4.1 Proof of part (i) of Theorem 6
$\left|\sqrt{\left|\|{\boldsymbol{\theta}}\|_{2}^{2}+2\sigma\boldsymbol{\theta}^{\top}{\boldsymbol{\varepsilon}}\right|}-\|{\boldsymbol{\theta}}\|_{2}\right|\leq\frac{2\sigma|\boldsymbol{\theta}^{\top}{\boldsymbol{\varepsilon}}|}{\|\boldsymbol{\theta}\|_{2}}.$ (161)
@@
$\left|\sum_{i=1}^{d}\Big(\varepsilon_{i}^{2}-\frac{\hat{\sigma}_{D}^{2}}{\sigma^{2}}\sigma_{i}^{2}\Big)\right|\leq\left|\sum_{i=1}^{d}(\varepsilon_{i}^{2}-\sigma_{i}^{2})\right|+\left|\sum_{i=1}^{d}\sigma_{i}^{2}\left(1-\frac{\hat{\sigma}_{D}^{2}}{\sigma^{2}}\right)\right|.$ (163)
Assumption 1 implies that $\sum_{i=1}^{d}\sigma_{i}^{2}\leq d$. Therefore, we can use Theorem 5 to get
@@ 6.4.2 Proof of parts (ii) and (iii) of Theorem 6
Note that since $\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}\leq d$ it follows from Theorem 4 that ${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}(\hat{\sigma}_{S}^{2})\lesssim\sigma^{2}$. Moreover, Assumption 1 implies that $\sum_{i\in\mathcal{S}}\sigma_{i}^{2}\leq s$. It follows that
@@
where (a) follows from Assumption 1 and Theorem 4 while (b) is due to the second inequality in Proposition 2(ii), the assumption that $s\leq\|{\boldsymbol{\Sigma}}\|_{F}$, the inequality $\alpha_{s}\lesssim\tau^{2}e^{-\tau^{2}/2}$ and the fact that $\tau=3\sqrt{\log(1+\|{\boldsymbol{\Sigma}}\|_{F}^{2}/s^{2})}$.
@@
$(s+\|{\boldsymbol{\Sigma}}\|_{F})\tau^{3}e^{-\tau^{2}/4}\lesssim s\log(1+\|{\boldsymbol{\Sigma}}\|_{F}^{2}/s^{2})=\psi^{*}(s,{\boldsymbol{\Sigma}}).$ (182)
@@
$\mathcal{A}:=\left\{\frac{|\hat{\sigma}_{S}^{2}-\sigma^{2}|}{\sigma^{2}}\leq\delta:=\frac{c(s+\|{\boldsymbol{\Sigma}}\|_{F})}{2\eta d}\right\},$ (189)
@@
$\int_{a}^{b}x^{2}e^{-x^{2}/2}dx\leq b^{2}e^{-a^{2}/2}(b-a).$ (194)
Consequently, since $\sigma_{i}^{2}\leq 1$ by Assumption 1, we have
${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left(\Big|\sum_{i\not\in\mathcal{S}}\varepsilon_{i}^{2}\left(\mathbbm{1}\left(|\varepsilon_{i}|>\frac{\hat{\sigma}_{S}}{\sigma}\sigma_{i}\tau\right)-\mathbbm{1}\left(|\varepsilon_{i}|>\sigma_{i}\tau\right)\right)\Big|\mathbbm{1}(\mathcal{A})\right)\lesssim\frac{s+\|{\boldsymbol{\Sigma}}\|_{F}}{\eta}\tau^{3}e^{-\tau^{2}/4}.$ (195)
@@
${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\left\{\left|\sum_{i\not\in\mathcal{S}}\varepsilon_{i}^{2}\left(\mathbbm{1}\left(|\varepsilon_{i}|>\frac{\hat{\sigma}_{S}}{\sigma}\sigma_{i}\tau\right)-\mathbbm{1}\left(|\varepsilon_{i}|>\sigma_{i}\tau\right)\right)\right|\geq\frac{C(s+\|{\boldsymbol{\Sigma}}\|_{F})}{\eta^{2}}\tau^{3}e^{-\tau^{2}/4}\right\}\cap\mathcal{A}\right]\leq\eta/2,$ (196)
@@
where, by the Schur complement (see, e.g., [11, (0.7.3.1)]),
@@
${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\frac{\hat{\sigma}_{S}(t)}{\sigma}\wedge 1\leq\frac{|\varepsilon_{i}|}{\sigma_{i}\tau}\leq\frac{\hat{\sigma}_{S}(t)}{\sigma}\vee 1\right]\lesssim\sqrt{({\boldsymbol{\Sigma}}^{-1})_{ii}}\frac{s+\|{\boldsymbol{\Sigma}}\|_{F}}{d}e^{-\frac{\tau^{2}}{16}}+\left(\frac{s+\|{\boldsymbol{\Sigma}}\|_{F}}{d}\right)^{2}.$ (199)
@@
$\hat{\sigma}_{-i}^{2}(t):=\frac{t}{F^{-1}(\hat{F}_{-i}(t))},\qquad\delta_{-i}(t):=\left|\frac{\hat{\sigma}_{-i}^{2}(t)}{\sigma^{2}}-1\right|,$ (200)
@@
$\left|\frac{\hat{\sigma}_{-i}^{2}(t)}{\sigma^{2}}-\frac{\hat{\sigma}_{S}^{2}(t)}{\sigma^{2}}\right|\leq c/d,$ (202)
@@
$\iota_{i}^{2}=\frac{1}{\sigma_{i}^{2}({\boldsymbol{\Sigma}}^{-1})_{ii}}\in[0,1],\quad\mu_{i}({\boldsymbol{\varepsilon}}_{-i})\sim\mathcal{N}(0,1-\iota_{i}^{2}).$ (206)
@@
${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[|\varepsilon_{i}|/\sigma_{i}\in[l,u],E_{1}\right]\leq\penalty 10000\$${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\mathbbm{1}(E_{1})\sqrt{\frac{2}{\pi}}\frac{L}{\iota_{i}}\exp\left(-\frac{(l-|\mu_{i}({\boldsymbol{\varepsilon}}_{-i})|)_{+}^{2}}{2\iota_{i}^{2}}\right)\right],$ (207)
@@
${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[|X|\in[a,b]\right]$ $=\int\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(x-\mu)^{2}}{2\sigma^{2}}}\mathbbm{1}(a\leq|x|\leq b)dx$ (208)
@@
$\phantom{{}={}}{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\mathbbm{1}(E_{1})L\exp\left(-\frac{(l-|\mu_{i}({\boldsymbol{\varepsilon}}_{-i})|)_{+}^{2}}{2\iota_{i}^{2}}\right)\right]$ (211)
@@
${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[L^{2}\mathbbm{1}(E_{1})\right]\lesssim\left(\frac{s+\|{\boldsymbol{\Sigma}}\|_{F}}{d}\right)^{2}.$ (214)
@@
${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\frac{\hat{\sigma}_{S}^{2}(t)}{\sigma^{2}}\wedge 1\leq\frac{\varepsilon_{i}^{2}}{\sigma_{i}^{2}\tau^{2}}\leq\frac{\hat{\sigma}_{S}^{2}(t)}{\sigma^{2}}\vee 1,E_{1}\right]\lesssim\sigma_{i}\sqrt{({\boldsymbol{\Sigma}}^{-1})_{ii}}\frac{s+\|{\boldsymbol{\Sigma}}\|_{F}}{d}\exp\left(-\frac{\tau^{2}}{16}\right).$ (215)
@@
${\rm E}_{1}=\{\hat{t}\in\mathcal{T}\},\quad{\rm E}_{2}=\left\{\hat{\sigma}_{S}^{2}\leq\frac{3}{2}\sigma^{2}\right\},$ (216)